# SynFlowNet: Design of Diverse and Novel Molecules with Synthesis Constraints

### v2

@@ 4.5 Scaling to the entire Enamine set
In this section, we study the ability of SynFlowNet to scale to large sets of building blocks using the method proposed in Section 3.3 as opposed to using a learnable embedding for each building block. First, we compare the performance of our models using the softmax action-selection mechanism and the fingerprints-based softmax on the full Enamine set of $221,181$ building blocks and randomly sampled subsets containing $75\%$, $50\%$, $25\%$, $10\%$, and $3\%$ of the full set. As shown in Figure 7A, increasing the number of building blocks in general negatively affects the quality of the sampled molecules in terms of reward. However, the model with fingerprints consistently outperforms the former and makes the reward degradation effect almost negligible. Next, we study whether models are able to use the blocks they are given. As shown in Figure 7B, the model with fingerprints consistently uses more unique building blocks than the baseline model. Notably, the baseline model is only able to use a small subset of building blocks to produce high-quality samples ($R>0.8$). In contrast, fingerprints enable the use of a 5-fold larger set of blocks for the same quality level, thus significantly surpassing the baseline samples in terms of diversity. Finally, we measure the average time of the forward pass and average time of forward pass with auxiliary RDKit operations like template matching. Figure 7C shows that increasing the number of building blocks slows down training, and using fingerprints introduces additional overhead.
As shown in Figure 7D, fingerprints enable more efficient learning on a large building block set. Indeed, instead of memorising all available BBs, the model learns to navigate in the space of embeddings and to maximise dot product with the relevant building blocks. To illustrate this, we clustered all available building blocks based on Tanimoto similarity. Next, we studied how both models explore new building blocks within and across these clusters. As shown in Figure 7E-F, the former model tends to explore more clusters and less unique building blocks within clusters. The model with fingerprints, on the contrary, is able to focus on clusters that maximise the reward.
@@ 4.6 Guiding SynFlowNet with experimental fragment screens
One advantage of our framework is that building block sets can be used to specialize the model for a particular target. Fragment-based Drug Discovery (FBDD) (Thomas et al., 2019) is a major strategy to increase the efficiency of drug discovery campaigns. In FBDD, instead of screening large chemical spaces, a low number of small compounds (called fragments) are screened and experimentally validated to bind, and are then linked in silico to make full molecules.
@@ Acknowledgements
We thank Enamine for providing the building blocks and for useful conversations. We thank Daniel Cutting, Austin Tripp and Jeff Guo for discussions regarding baselines. Miruna Cretu is supported by the SynTech CDT at the University of Cambridge and part of the work was completed during an internship at Valence Labs.
@@ A.1 Reward functions
Finally, we wish for users to rapidly be able to adapt SynFlowNet to learn binding for their target of interest without having to retrain a new proxy or relying on slow docking simulations. We accomplish this using the new GPU-accelerated Vina-GPU 2.1 docking algorithm (Tang et al., 2023; Alhossary et al., 2015). For our experiment, we use the PDB:6W63 (Mesecar et al., 2020) structure for the SARS-CoV-2 main protease (Mpro) target and PDB:2XJX (Murray et al., 2010) for Heat Shock Protein 90 (HSP90) target. Receptors are prepared for docking with prepare_receptor4.py and the center of the docking is defined as the center of mass for the ligand with a size of 25 $\mathrm{\SIUnitSymbolAngstrom}$ in accordance with previous work (Buttenschoen et al., 2024).
A.2 Chemical reactions
Figure A.1: Example of reaction templates (Hartenfeller et al., 2012). The templates act as rules, which match any molecule that left side (before the arrow) as a subgraph. The matching part is then transformed into the right hand side of the rule. “R” represents any group, “A” represents an aromatic atom. Note the implicit reagent used in reaction #2.
@@ A.2 Chemical reactions
When running experiments with a REINFORCE backward policy (which requires to sample trajectories backward on-policy), we enforced additional masking that forward reactions are sampled if and only if they are reversible (i.e. when applying the template backward on the product, the same reactants are obtained as the ones utilised by the template in the forward direction).
Note that reaction templates can be uni-molecular (employing one reactant, Fig. A.1 #2) or bi-molecular (employing two reactants, Fig. A.1 #1).
@@ A.5 SynFlowNet scaling
For Section 4.5, to compute clusters of building blocks, we used BitBIRCH (Jung et al., 2024), a recent adaptation of Balanced Iterative Reducing and Clustering using Hierarchies (BIRCH) algorithm (Zhang et al., 1996), recently proposed for efficient clustering of large molecular libraries.
@@ B.1 Comparison to baselines
Table A.2: Comparison to baselines. Results obtained by averaging over 1000 random molecules sampled from the trained models. Standard errors obtained from training using 3 seeds. Due to high computational cost, AiZynthFinder scores are computed over 100 random samples.
@@
B.2 Example Trajectories
@@
B.3 Curated building blocks sets given target data
@@
For every experimentally validated fragment, we selected the top 100 closest building block molecules from the Enamine library based on molecular similarity. After removing duplicate SMILES entries, this process resulted in thousands of curated building blocks for each target.

### v3

@@ 4.5 Scaling to the entire Enamine set
In this section, we study the ability of SynFlowNet to scale to large sets of building blocks using the method proposed in Section 3.3 as opposed to using a learnable embedding for each building block. First, we compare the performance of our models using the softmax action-selection mechanism and the fingerprints-based softmax on the full Enamine set of $221,181$ building blocks and randomly sampled subsets containing $75\%$, $50\%$, $25\%$, $10\%$, and $3\%$ of the full set. As shown in Figure 6A, increasing the number of building blocks while using the same number of training steps negatively affects the reward for the model without fingerprints. However, the model with fingerprints consistently outperforms the former and makes the reward degradation effect almost negligible. Next, we study whether models are able to use the blocks they are given. As shown in Figures 6B, the model with fingerprints consistently uses more unique building blocks than the baseline model. Notably, the baseline model is only able to use a small subset of building blocks to produce high-quality samples ($R>0.8$). In contrast, fingerprints enable the use of a 5-fold larger set of blocks for the same quality level, thus significantly surpassing the baseline samples in terms of diversity. Finally, we measure the average time of the forward pass and average time of forward pass with auxiliary RDKit operations like template matching in Figure 6C.
As shown in Figure 6D, fingerprints enable more efficient learning on a large building block set. Indeed, instead of memorising all available BBs, the model learns to navigate in the space of embeddings and to maximise dot product with the relevant building blocks. To illustrate this, we clustered all available building blocks based on Tanimoto similarity. Next, we studied how both models explore new building blocks within and across these clusters. As shown in Figures 6E-F, the baseline model tends to explore more clusters while discovering low-reward molecules, thus becoming less efficient. The model with fingerprints, on the contrary, is able to focus on clusters that maximise the reward, leveraging the structure of the chemical space embedded in fingerprints.
@@ 4.6 Guiding SynFlowNet with experimental fragment screens
One advantage of our framework is that building block sets can be used to specialize the model for a particular target. Fragment-based Drug Discovery (FBDD) (Thomas et al., 2019) is a major strategy to increase the efficiency of drug discovery campaigns. In FBDD, instead of screening large chemical spaces, a relatively low number of small compounds (called fragments) are screened and experimentally validated to bind, and are then linked or merged in silico to make full molecules.
@@ Acknowledgements
We thank Enamine for providing the building blocks and for useful conversations. We thank Daniel Cutting, Austin Tripp and Jeff Guo for discussions regarding baselines. We thank Doo-Hyun Kwon (Recursion) for running the SpaceLight software. Miruna Cretu is supported by the SynTech CDT at the University of Cambridge and part of the work was completed during an internship at Valence Labs.
@@ A.1 Reward functions
Finally, we wish for users to rapidly be able to adapt SynFlowNet to learn binding for their target of interest without having to retrain a new proxy or relying on slow docking simulations. We accomplish this using the new GPU-accelerated Vina-GPU 2.1 docking algorithm (Tang et al., 2023; Alhossary et al., 2015). For our experiments, we use the PDB:6W63 (Mesecar et al., 2020) structure for the SARS-CoV-2 main protease (Mpro) target, PDB:2XJX (Murray et al., 2010) for Heat Shock Protein 90 (HSP90) target and PDB:8AZR for KRAS (Kim et al., 2023). Receptors are prepared for docking with prepare_receptor4.py and the center of the docking is defined as the center of mass for the ligand with a size of 25 $\mathrm{\SIUnitSymbolAngstrom}$ in accordance with previous work (Buttenschoen et al., 2024). In order to prevent the model from optimising for large molecules, we add a reward penalty of $-0.4$ for molecules with a number of heavy atoms larger than that of the reference ligand plus an allowance of 8 additional heavy atoms. The Vina scores are scaled between 0 and 1 according to:
$R=\frac{affinities+reward\_scale\_min}{reward\_scale\_min+reward\_scale\_max}-1,$ (4)
where $reward\_scale\_min$ and $reward\_scale\_max$ are tunable and set to -1.0 and respectively -10.0 by default.
A.2 Data
We use commercially available building blocks (BBs) from Enamine , which are small fragments of molecules prepared in bulk to be readily synthesised into candidate molecules. Reaction templates are obtained from two publicly available template libraries (Button et al., 2019; M et al., 2012). After preprocessing, we obtain a total of 105 templates: 13 uni- and 92 bi-molecular reactions respectively (see App. A.2.1). We also use 12 Enamine REAL reactions in a small number of experiments (see Section 4.6 and App. A.2.1).
A.2.1 Chemical reactions
Figure A.1: Example of reaction templates (M et al., 2012). The templates act as rules, which match any molecule that left side (before the arrow) as a subgraph. The matching part is then transformed into the right hand side of the rule. “R” represents any group, “A” represents an aromatic atom. Note the implicit reagent used in reaction #2.
@@ A.2.1 Chemical reactions
When running experiments with a REINFORCE backward policy (which requires to sample trajectories backward on-policy), we enforced additional masking that forward reactions are sampled if and only if they are reversible (i.e. when applying the template backward on the product, the same reactants are obtained as the ones utilised by the template in the forward direction). Note that reaction templates can be uni-molecular (employing one reactant, Fig. A.1 #2) or bi-molecular (employing two reactants, Fig. A.1 #1).
@@ A.5 SynFlowNet scaling
Here, we provide technical details of the experiments described in Section 4.5. For each building block subset, we trained 3 models using fingerprints, and 3 models without fingerprints, which amounts to 36 models in total. All models had exactly the same hyper-parameter configuration with SeH proxy as reward, 8000 training steps, batch size 8 and $\beta=32$. Other parameters were the same as in Table LABEL:tab:hyperparam.
For evaluation, we sampled 5000 trajectories with each model and considered only trajectories with valid final molecules. In Figure 6C, we report average runtime per batch over the training process. model.forward() corresponds to the average time required to perform the matrix operations (PyTorch back-end, single GPU NVIDIA H100) during the forward pass of the model and aims to highlight the performance overhead introduced by the dot product with molecular fingerprints. model.forward() + RDKit additionally accounts for the costly template matching operations ran on CPU after the model’s forward pass (to shortlist the set of allowed building blocks for the chosen reaction). These values demonstrate the effect of the increased building block set. To characterize the structural diversity of the molecules generated by different models, we additionally provide the average pairwise Tanimoto dissimilarity of the molecules sampled by different models (using different building block subsets) in Figure A.2.
For Figures 6D-F, the entire building block set was used. To compute clusters of building blocks for Figures 6E-F, we used BitBIRCH (Jung et al., 2024), a recent adaptation of Balanced Iterative Reducing and Clustering using Hierarchies (BIRCH) algorithm (Zhang et al., 1996), recently proposed for efficient clustering of large molecular libraries.
Figure A.2: Diversity of the molecules produced by the baseline and fingerprint models trained with different building block subsets.
@@ Appendix B Extended Results
Figure A.3: Example of synthesis pathway generated by SynFlowNet. The reward here is the sEH binding affinity proxy. Right shows the Vina docked pose and molecule metrics. More examples are shown in Figure A.7.
@@ B.1 Comparison to baselines
Table A.2: Comparison to baselines. Results obtained by averaging over 1000 random molecules sampled from the trained models. Standard errors obtained from training using 3 seeds. SynFlowNet ChEMBL refers to a model trained with building blocks derived from ChEMBL molecules. Due to high computational cost, AiZynthFinder scores are computed over 100 random samples.
Method KRAS Vina ($\uparrow$) Diversity ($\uparrow$) SA ($\downarrow$) QED ($\uparrow$) Mol. weight ChEMBL Similarity ($\downarrow$)
REINVENT $0.76\pm 0.03$ $0.85\pm 0.01$ $2.51\pm 0.05$ $0.53\pm 0.03$ $359.52\pm 6.51$ $0.57\pm 0.01$
FragGFN $0.86\pm 0.06$ $0.62\pm 0.11$ $3.54\pm 0.02$ $0.43\pm 0.03$ $348.89\pm 34.42$ $0.42\pm 0.04$
Soft Q-Learning $0.66\pm 0.01$ $0.88\pm 0.02$ $3.10\pm 0.05$ $0.61\pm 0.01$ $370.97\pm 12.76$ $0.44\pm 0.02$
SynFlowNet ($L=3$) $0.73\pm 0.09$ $0.87\pm 0.01$ $3.05\pm 0.10$ $0.56\pm 0.08$ $389.91\pm 23.27$ $0.42\pm 0.01$
Table A.3: KRAS binding optimization with GPU-accelerated Vina docking. Results obtained by averaging over 1000 random molecules sampled from the trained models. Standard errors obtained from training using 3 seeds. Vina scores are scaled between 0 and 1 using Eq. 4.
@@
We further compare SynFlowNet’s performance to SynNet’s (Gao et al., 2022b), which models the generation of synthetic trees containing multi-step synthesis pathways as a Markov decision process. The approach differs from SynFlowNet and Gottipati et al. (2020a); Horwood & Noutahi (2020) in that it relies on a trained model optimized using a genetic algorithm, instead learning the target distribution solely from the reward function, based on an RL objective. In Table A.4 we report top-100 performance on the optimization tasks performed in SynNet.
Table A.4: Comparison to SynNet. We report average of top-100 scores. Results for SynNet are reported from Gao et al. (2022b).
[…]
SynNet 0.947 0.815 0.998 0.719
SynFlowNet 0.947 0.862 0.999 0.710
Novelty
For a more strict evaluation of the novelty of SynFlowNet-proposed designs, we investigate whether SynFlowNet can still generate low-similarity molecules to ChEMBL when relying on building blocks derived from ChEMBL actives. In this way, we reach comparable vocabularies between REINVENT’s prior and SynFlowNet’s action space. For this, we select 70 0007000070\,00070 000 random ChEMBL molecules from Zdrazil et al. (2023) and run AiZynthFinder retrosynthesis (Genheden et al., 2020) to decompose the molecules into building blocks. This results in $8527$ unique building blocks which were used to train SynFlowNet. We report comparative results from this experiment in Figure A.5 and Table A.3 and note that SynFlowNet preserves high novelty compared to REINVENT, which emphasizes the advantages of the framework.
Figure A.5: SynFlowNet trained with ChEMBL-derived building blocks remains competitive against other popular models from the literature. SynFlowNet achieves the best novelty/reward and diversity/reward trade-offs.
B.1.1 Sample efficiency
[…]
Table A.5: Sample efficiency results. We report mean and standard deviation of the AUC of top-10 average scores versus the number of oracle calls from 5 independent runs. Results for all other methods are reported from Gao et al. (2022a).
Task Method AUC
[…]
SynNet $0.789\pm 0.032$
FragGFN $0.651\pm 0.026$
SynFlowNet $0.691\pm 0.034$
DRD2 REINVENT $0.943\pm 0.005$
SynNet $0.969\pm 0.004$
FragGFN $0.590\pm 0.070$
SynFlowNet $0.885\pm 0.027$
@@ B.1.1 Sample efficiency
B.2 Outlook
While SynFlowNet shows promising results for generating high-quality synthesizable molecules, there remain challenges to be addressed in future work. First, SynFlowNet does not explore non-linear synthetic pathways, i.e. considering intermediates in the DAG as second reactants. At the moment the choice of second reactants is limited to the building blocks library. SynFlowNet also does not handle reactions with more than 2 reactants. We also emphasize that reaction selectivity is not accounted for, and that including reaction feasibility in the design process would strengthen real-world synthesizability results. Another inherent limitation of the reaction templates employed in this work is the lack of consideration of reaction conditions and stereochemistry. We envision that the backward policy in SynFlowNet can be an innovative way of accounting for additional constraints in the MDP design, and our preliminary results in Section 3.2 on training the backward policy with a separate reward from the forward policy support this. One could further bake in constraints such as synthesis cost, reaction yield and selectivity in the backward reward $R_{B}$. Additionally, a more careful curation of the reaction set is encouraged in future work, which would enhance the coverage of the chemical space via diverse transformations: cyclizations, side-chain extensions, functional group interconversions etc. Lastly, the rediscovery rate of SynFlowNet is not perfect, meaning that certain regions of the chemical space may remain inaccessible to the model under the current datasets.
B.3 Example Trajectories
@@
B.4 Curated building block sets given target data
@@
For every experimentally validated fragment, we selected the top 100 closest building block molecules from the Enamine library based on molecular similarity. After removing duplicate SMILES entries, this process usually resulted in thousands of curated building blocks for each target.